OpenStudy (anonymous):

What is the equation of the line perpendicular to the line y =4/9 x – 2 that passes through the point (4,3)?

OpenStudy (triciaal):

perpendicular lines the product of the slopes = -1 line 1 slope = m slope of perpendicular = -1/m what is the slope of the given line?

OpenStudy (triciaal):

line y =4/9 x – 2 equate with y = mx + b what is m?

OpenStudy (triciaal):

are you going to answer? do you still need help or no?

OpenStudy (anonymous):

sorry i got the answer of y=2/7x-12